# Differentiable Structure Learning for General Binary Data

### v1

@@ C.4 Procedure for recovering causal graph and parameters
3    $\mathrm{PA}(\pi(j))\leftarrow\mathrm{PA}\cup\{\pi(i)\}$
@@ C.5 Structural equation model
where $f=(f_{j})_{j=1}^{p}$ is a collection of functions $f_{j}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{p+1}\rightarrow\mathbb{R}$, here $N=(N_{1},\ldots,N_{p})$ is a vector of independent noises with distribution $P(N)$, and $\mathrm{PA}(j)$ denotes the set of parents of node $j$. Here, $\partial_{k}f_{j}$ denotes the partial derivative of $f_{j}$ w.r.t. $X_{k}$, which is identically zero when $f_{j}$ is independent of $X_{k}$, i.e. $f_{j}(X,N_{j})=f_{j}(X_{\mathrm{PA}(j)},N_{j})$.
The graphical structure induced by the SEM, assumed to be a DAG, will be represented by the following $p\times p$ weighted adjacency matrix $B$:
@@
C.6 Faithfulness, sparest Markov representation, Markov equivalence class
@@ C.8 Theoretical justification for previous works
$\begin{split}\mathbf{H}_{j}=&(\underbrace{h^{j,0}}_{\text{constant}},\underbrace{h^{j,1},\ldots,h^{j,p}}_{\text{first order }})^{\top}\in\mathbb{R}^{p+1}\qquad\mathbf{H}=(\mathbf{H}_{1},\ldots,\mathbf{H}_{p})\in\mathbb{R}^{(p+1)\times p}\end{split}$ (87)
where $h^{j,0}$ is the intercept and $h^{j,1},\dots,h^{j,p}$ are the first‐order coefficients.
@@ 20 $\mathbf{X}_{\pi(j)}\sim\mathrm{Bernoulli}(q)$
Implementation
For each dataset, we applied several structural learning algorithms, including fast greedy equivalence search (FGES [Ramsey et al., 2017]), constraint-based methods (PC [Spirtes et al., 2000]), DAGMA [Bello et al., 2022], NOTEARS [Zheng et al., 2020]. The implementation details are provided in the following paragraph. After running the algorithms, a post-processing threshold of $0.3$ was applied to the estimated $B_{\mathrm{est}}$ to prune small values, following the same procedure in Zheng et al. [2018], Bello et al. [2022].
@@
  • NOTEARS‐MLP [Zheng et al., 2020] is a continuous differentiable DAG‐learning method that employs a least‐squares loss with $\ell_{1}$ regularization, solved by augmented langrage method. Its Python implementation is available at https://github.com/xunzheng/notears. In our variant, we insert a sigmoid activation $\sigma(x)=1/(1+\exp(-x))$ on the final layer and replace the original loss with the cross‐entropy (logistic) loss to accommodate binary data. After estimating the weighted adjacency matrix $W_{\mathrm{est}}$ via NOTEARS‐MLP, we prune all entries below a threshold of $0.3$, compute a topological ordering of the resulting graph, and then apply Algorithm 2 with first and second order to obtain the final structure. Finally, we remove any remaining edges whose weight does not exceed $1.0$ to eliminate spurious connections. We name this method as Notears-mlp-reg. These heuristic methods are applied to larger graphs with $d\in\{10,20,30,40\}$.
  • DAGMA [Bello et al., 2022] is a continuous DAG‐learning algorithm that achieves improved accuracy and faster computation, with barrier methods. Its implementation can be found at https://github.com/kevinsbello/dagma. To highlight that the original DAGMA only models first‐order interactions, we refer to it as Dagma-1st. By solving (13) while incorporating all higher‐order interactions, we arrive at our extended method, denoted Dagma-ho. For small graphs, we implement the full formulation (13) including all higher-order interactions; hence, Dagma-ho is applied for $d\in\{5,6,7,8,9\}$.
@@
Theorem 3 indicates that one should ideally choose small values of $\lambda$ and $\delta$ for the quasi-MCP penalty. In practice, however, achieving the global optimum of (13) is infeasible, and if $\lambda$ and $\delta$ are too small the penalty becomes ineffective and the algorithm may fail to recover the sparsest solution. To mitigate this, we adopt the continuation strategy of Deng et al. [2024]: start with relatively large $\lambda$ and $\delta$, solve (13) via Dagma-ho to obtain an initial estimate $\mathbf{H}_{\mathrm{est}}$, then iteratively shrink $\lambda$ and $\delta$ by a factor $\gamma<1$, using the previous estimate as the warm start for the next run of Dagma-ho. We terminate when the negative log-likelihood $s(\mathbf{H}_{\mathrm{est}};{\mathbf{X}})$ ceases to decrease. Empirically, $\gamma=0.5$, $\lambda=0.05$, and $\delta=0.2$ perform well in our experiments.
Equipment
The experiments are conducted in the following CPU architectures
  • Intel Broadwell—28 cores @ 2.4 GHz with 64 GB memory per node
  • Intel Skylake—40 cores @ 2.4 GHz with 96 GB memory per node
D.2 Metrics
@@ E.1 Small graphs
Figure 3: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Column: $k=\{1,2\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{5,6,7,8,9\}$. DAGMA [Bello et al., 2022] is renamed as “DAGMA-1st”, to emphasize only linear term is used. Error bars denote the standard error computed over 10 replications.
@@ E.2 Large graphs
Figure 4: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Data are generated using extended feature map $\Phi^{\text{1st+pth}}$ in (99). Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. Notears-mlp-reg is our two stage approach. DAGMA [Bello et al., 2022] is renamed as “DAGMA-1st”, to emphasize only linear term is used.
Figure 5: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Data are generated using extended feature map $\Phi^{\text{2nd}}$ in (99). Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. Notears-mlp-reg is our two stage approach. DAGMA [Bello et al., 2022] is renamed as “DAGMA-1st”, to emphasize only linear term is used.
Figure 6: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Data are generated using extended feature map $\Phi^{\text{pth}}$ in (99). Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. Notears-mlp-reg is our two stage approach. DAGMA [Bello et al., 2022] is renamed as “DAGMA-1st”, to emphasize only linear term is used.

### v2

@@ C.4 Procedure for recovering causal graph and parameters
3    $\mathrm{PA}(\pi(j))\leftarrow\mathrm{PA}(\pi(j))\cup\{\pi(i)\}$
@@ C.5 Structural equation model
where $f=(f_{j})_{j=1}^{p}$ is a collection of functions $f_{j}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{p+1}\rightarrow\mathbb{R}$, here $N=(N_{1},\ldots,N_{p})$ is a vector of independent noises with distribution $P(N)$, and $\mathrm{PA}(j)$ denotes the set of parents of node $j$. Here, $\partial_{k}f_{j}$ denotes the partial derivative of $f_{j}$ w.r.t. $X_{k}$, which is identically zero when $f_{j}$ is independent of $X_{k}$, i.e. $f_{j}(X,N_{j})=f_{j}(X_{\mathrm{PA}(j)},N_{j})$. The graphical structure induced by the SEM, assumed to be a DAG, will be represented by the following $p\times p$ weighted adjacency matrix $B$:
@@
C.6 Faithfulness, Sparsest Markov representation, Markov equivalence class
@@ C.8 Theoretical justification for previous works
$\begin{split}\mathbf{H}_{j}=&(\underbrace{h^{0,j}}_{\text{constant}},\underbrace{h^{1,j},\ldots,h^{p,j}}_{\text{first order }})^{\top}\in\mathbb{R}^{p+1}\qquad\mathbf{H}=(\mathbf{H}_{1},\ldots,\mathbf{H}_{p})\in\mathbb{R}^{(p+1)\times p}\end{split}$ (92)
where $h^{0,j}$ is the intercept and $h^{1,j},\dots,h^{p,j}$ are the first‐order coefficients.
@@ 20 $\mathbf{X}_{\pi(j)}\sim\mathrm{Bernoulli}(q)$
D.2 Implementation
For each dataset, we applied several structural learning algorithms, including fast greedy equivalence search (FGES [Ramsey et al., 2017]), constraint-based methods (PC [Spirtes et al., 2000]), DAGMA [Bello et al., 2022], NOTEARS [Zheng et al., 2020]. The implementation details are provided in the following paragraph. After running the algorithms, a post-processing threshold of $0.3$ was applied to the estimated $B_{\mathrm{est}}$ to prune small values, following the same procedure in [Zheng et al., 2018, Bello et al., 2022].
@@ D.2 Implementation
  • NOTEARS‐MLP [Zheng et al., 2020] is a continuous differentiable DAG‐learning method that employs a least‐squares loss with $\ell_{1}$ regularization, solved by augmented lagrangian method. Its Python implementation is available at https://github.com/xunzheng/notears.
  • DAGMA [Bello et al., 2022] is a continuous DAG‐learning algorithm that achieves improved accuracy and faster computation, with barrier methods. Its implementation can be found at https://github.com/kevinsbello/dagma. We use the default parameters from original implementation.
  • BiNOTEARS builds on DAGMA and NOTEARS-MLP, with the choice of solver tailored to problem size. For small graphs ($d\in\{5,6,7,8,9\}$), we optimize the full formulation (13)—including all higher-order interaction terms—using the DAGMA framework. For larger graphs ($d\in\{10,20,30,40\}$), we adopt the NOTEARS-MLP framework. The original implementation optimizes an $\ell_{2}$ loss with an $\ell_{1}$ penalty; in our variant, we insert a sigmoid activation $\sigma(x)=1/(1+\exp(-x))$ on the final layer and replace the squared loss with the cross-entropy (logistic) loss to accommodate binary data. After estimating the weighted adjacency $B_{\mathrm{est}}$ via NOTEARS-MLP, we prune entries below $0.3$, compute a topological ordering of the resulting graph, and then apply Algorithm 2 with first- and second-order terms to obtain the final structure. Finally, we remove any remaining edges whose weight does not exceed $1.0$ to eliminate spurious connections.
@@
Theorem 3 indicates that one should ideally choose small values of $\lambda$ and $\delta$ for the quasi-MCP penalty. In practice, however, achieving the global optimum of (13) is infeasible, and if $\lambda$ and $\delta$ are too small the penalty becomes ineffective and the algorithm may fail to recover the sparsest solution. To mitigate this, we adopt the continuation strategy of Deng et al. [2024]: start with relatively large $\lambda$ and $\delta$, solve (13) via BiNOTEARS to obtain an initial estimate $\mathbf{H}_{\mathrm{est}}$, then iteratively shrink $\lambda$ and $\delta$ by a factor $\gamma<1$, using the previous estimate as the warm start for the next run of BiNOTEARS. We terminate when the negative log-likelihood $s(\mathbf{H}_{\mathrm{est}};{\mathbf{X}})$ ceases to decrease. Empirically, $\gamma=0.5$, $\lambda=0.05$, and $\delta=0.2$ perform well in our experiments.
D.3 Metrics
@@ E.1 Small graphs
Figure 3: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Column: $k=\{1,2\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{5,6,7,8,9\}$. Error bars denote the standard error computed over 10 replications.
@@ E.2 Large graphs
Figure 4: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Data are generated using extended feature map $\Phi^{\text{1st+pth}}$ in (104). Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. BiNOTEARS is our two stage approach.
Figure 5: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Data are generated using extended feature map $\Phi^{\text{2nd}}$ in (104). Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. BiNOTEARS is our two stage approach.
Figure 6: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Data are generated using extended feature map $\Phi^{\text{pth}}$ in (104). Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. BiNOTEARS is our two stage approach.